OpenStudy (anonymous):

What is the sum of the measures of the interior angles of a regular polygon where a single exterior angle measures 30°?

OpenStudy (anonymous):

Exterior angle = 30 Sum of all exterior angles = 360 Number of sides of polygon = n = 360/30 = 12 Sum of all interior angles = 180(n-2) = 180(12-2) = 1800 degrees